ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

1. Logic

1991 MS Classification: 03,04 Zeman, Martin, Humboldt Uni, Germany ◇ at Mahlo Cardinals Let κ be Mahlo, ε be regular uncountable and E_{ε} be the set of all cardinals less than κ of cofinality ε. If ◇_{κ}(E_{ε}) fails, then there are stationarily many β < κ such that o(β) ≥ ε in the core model. References: 1. Jensen, R. B. : On ◇ at Mahlo cardinals, handwritten notes, Oberwolfach, 1991 2. Zeman, M. : ◇ at Mahlo cardinals, preprint